OpenStudy (anonymous):

hoose the equation of the line passing through the point (-2, 4) and perpendicular to y = -one halfx + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x

OpenStudy (blurbendy):

do you know the point slope formula?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Since the lines are perpendicular, the product of the two slopes is -1 i.e m1 * m2 = -1 u have m2 = -1/2

OpenStudy (anonymous):

so the uqation of ur new line is y = mx+b m = m1 that u ll calculate and u r give (x,y) = (-2,4) just plug in the eq and u ll have the answer